After finding the sample proportion , and the margin of Error E, express the confidence interval 0.4 < p < 0.64 in the form of ± E.
Answer #1

Solution :

Point estimate = = (Lower confidence interval + Upper confidence interval ) / 2

Point estimate = = (0.4+0.64) / 2=0.52

Margin of error = E = Upper confidence interval - = 0.65-0.52=0.13

In the form of :

-+ E

0.52±0.1